# Order isomorphisms of sup-stable function spaces: continuous, Lipschitz, c-convex, and beyond

### v1

@@ 3.3 Continuous affine maps and the lower semicontinuous convex functions
As $G$ is bijective, $J^{-1}$ commutes with the max, and as the $p$-coefficient of $J^{-1}(Jf_{\alpha}+\epsilon)$ does not depend on $\epsilon$, we have that $J^{-1}(Jf_{\alpha}+\epsilon)=\langle p_{\alpha},\cdot\rangle+\lambda_{\alpha,\epsilon}$ for some $\lambda_{\alpha,\epsilon}$. We have furthermore that $\lambda_{\alpha,\epsilon}>\lambda_{\alpha}$ since $G_{2}(\_,\cdot)$ is bijective and increasing in $\lambda$. Hence, applying $J^{-1}$ to (40), we derive that
@@
We have just shown that $G$ maps intervals onto intervals. Arguing as in [AAM09, Remark 6, p668] by applying a result of affine geometry, extended in [IRS15, Corollary 3] to arbitrary vector spaces, $G$ is thus affine and bijective, i.e. we can fix a linear invertible $A:\mathscr{X}^{*}\to\mathscr{X}^{*}$, a linear form $C:\mathscr{X}^{*}\to\mathbb{R}$, and constants $g,b\in\mathscr{X}^{*}$, $d,\delta\in\mathbb{R}$ such that
@@
so, considering nets, we have that $A:\mathscr{X}^{*}\to\mathscr{X}^{*}$ is weak-* continuous as $g(x,\cdot)$ is continuous. Hence $A$ has a (weakly) continuous adjoint $A^{*}:\mathscr{X}\to\mathscr{X}$. Fix any convex l.s.c $f$. Then
@@
Conversely, take any $J$ satisfying (39). Such $J$ is clearly order preserving, and if we set $J^{1}f(x)=-\frac{1}{d}\langle b,A^{-1}(x-c)\rangle-\frac{\delta}{d}+\frac{1}{d}\cdot f(A^{-1}(x-c))$, it satisfies $J^{1}Jf=f$.
Applying 11, we know that, for $\mathcal{G}=\operatorname*{Rg}(B)$ for some $\bar{B}$ as in (10), the anti-involutions are only the composition of the reference one $\bar{B}$ with order isomorphisms. In this case, the reference is the Fenchel transform, and we thus recover [AAM09, Theorem 5] for $\mathscr{X}=\mathbb{R}^{d}$ as a special case and derive a straightforward extension to semi-convex functions.
@@
Let $\mathcal{G}$ be the space of proper convex lower semicontinuous functions over a reflexive Banach space $\mathscr{X}$ of dimension larger than two, assumed to be linearly isomorphic to its dual $\mathscr{X}^{*}$. Then the anti-involutions over $\mathcal{G}$ are of the form
@@
with $A\in\mathcal{L}(\mathscr{X})$ invertible, $c\in\mathscr{X}$, $\delta\in\mathbb{R}$ and $K^{-1}A^{-\top}KA=\operatorname*{Id}_{\mathscr{X}}$ where $K:\mathscr{X}\rightarrow\mathscr{X}^{*}$ is the duality operator and $A^{-\top}$ is the inverse of the dual of $A$.
@@ Acknowledgments.
The authors deeply thank Sebastián Tapia-García for the fruitful discussions and his many insights on topology while writing this article.
@@
$\forall f,g\in\mathscr{F}\text{ s.t.\ }f\leq g,\bar{F}(f)=\bar{F}(\min(f,g))=\max(\bar{F}(f),\bar{F}(g))\geq\bar{F}(g).$
Hence, as $\inf\nolimits^{\mathscr{F}}_{\alpha\in\mathscr{A}}f_{\alpha}\leq f_{\beta}$ for any $\beta\in\mathscr{A}$,
Since $\mathcal{G}$ is a complete subspace of ¯$\overline{\mathbb{R}}^{\mathscr{X}}$, $\sup_{\alpha\in\mathscr{A}}\bar{F}(f_{\alpha})\in\mathcal{G}$, so composing by $\bar{F}^{1}$ and using that $\bar{F}^{1}$ is order reversing, we derive that
$\bar{F}^{1}\bar{F}(\inf\nolimits^{\mathscr{F}}_{\alpha\in\mathscr{A}}f_{\alpha})\leq\bar{F}^{1}(\sup_{\alpha\in\mathscr{A}}\bar{F}(f_{\alpha}))\leq\bar{F}^{1}\bar{F}(f_{\beta}).$
$\inf\nolimits^{\mathscr{F}}_{\alpha\in\mathscr{A}}f_{\alpha}=\bar{F}^{1}\bar{F}(\inf_{\alpha\in\mathscr{A}}f_{\alpha})\leq\bar{F}^{1}(\sup_{\alpha\in\mathscr{A}}\bar{F}(f_{\alpha}))\leq\inf\nolimits^{\mathscr{F}}_{\beta\in\mathscr{A}}f_{\beta}.$
So $\inf\nolimits^{\mathscr{F}}_{\alpha\in\mathscr{A}}f_{\alpha}=\bar{F}^{1}\left(\sup_{\alpha\in\mathscr{A}}\bar{F}(f_{\alpha})\right)$, and, composing again by $\bar{F}$, we deduce $\bar{F}$ sends relative infima with respect to $\mathscr{F}$ to relative suprema of $\mathcal{G}$ (which coincide with ordinary suprema). Define now $\tilde{F}:\overline{\mathbb{R}}^{\mathscr{Y}}\rightarrow\overline{\mathbb{R}}^{\mathscr{X}}$ by, for any $f\in\overline{\mathbb{R}}^{\mathscr{Y}}$, $\tilde{F}(f):=\bar{F}(\inf\nolimits^{\mathscr{F}}f)\in\mathcal{G}$. Similarly, set $\tilde{F}^{1}:\overline{\mathbb{R}}^{\mathscr{X}}\rightarrow\overline{\mathbb{R}}^{\mathscr{Y}}$ as for any $g\in\overline{\mathbb{R}}^{\mathscr{X}}$, $\tilde{F}^{1}(g):=\bar{F}^{1}(\inf\nolimits^{\mathcal{G}}g)\in\mathscr{F}$.

### v2

@@ 3.3 Continuous affine maps and the lower semicontinuous convex functions
Set $f^{\prime}_{0}=Jf_{0}-Jf_{\alpha}$, $f^{\prime}_{1}=Jf_{1}-Jf_{\alpha}$, $F=\max(f^{\prime}_{0},f^{\prime}_{1})>0$. Denote by $G_{f}\subset\mathscr{X}\times\mathbb{R}$ the graph of a function $f:\mathscr{X}\to\mathbb{R}$. Let $\mathcal{C}:=G_{f^{\prime}_{0}}\cap G_{f^{\prime}_{1}}$. It is nonempty, as otherwise the hyperplanes are parallel, $G_{1}(p_{0})=G_{1}(p_{1})$, so $p_{0}=p_{1}$. Take $(x^{\prime}_{0},\lambda^{\prime}_{0})\in(G_{F}\cap G_{f^{\prime}_{0}})\backslash\mathcal{C}$ and $(x^{\prime}_{1},\lambda^{\prime}_{1})\in(G_{F}\cap G_{f^{\prime}_{1}})\backslash\mathcal{C}$, both sets being non-empty since $f^{\prime}_{0}$ is not parallel to $f^{\prime}_{1}$. We are down to a unidimensional reasoning in $P=(\{x^{\prime}_{0}\}+\operatorname*{span}(x^{\prime}_{1}-x^{\prime}_{0}))\times\mathbb{R}$. There clearly exists a unique $s\in(0,1)$ such that $x^{\prime}_{s}=(1-s)x^{\prime}_{0}+sx^{\prime}_{1}$ satisfies $(x^{\prime}_{s},F(x^{\prime}_{s}))\in\mathcal{C}$. If w.l.o.g. $F(x^{\prime}_{0})<F(x^{\prime}_{s})$, then since $\mathcal{C}$ is a vector space, $x(\cdot):\beta\in(-\infty,s]\mapsto(1-\beta)x^{\prime}_{0}+\beta x^{\prime}_{1}$ is such that $F(x(\beta))=f^{\prime}_{0}(x(\beta))$. Taking the limit $\beta\to-\infty$, by linearity of $f^{\prime}_{0}$, we obtain $F(x(\beta))\to-\infty$, while $F$ is lower bounded. Hence we must have $F(x_{s})\leq\min(F(x^{\prime}_{0}),F(x^{\prime}_{1}))$, and to estimate $\inf_{\mathscr{X}}F$, we can focus exclusively on $\mathcal{C}$. Take now $(x^{\prime}_{0},\lambda^{\prime}_{0})\in\mathcal{C}$ and $(x^{\prime}_{1},\lambda^{\prime}_{1})\in\mathcal{C}$ with $x^{\prime}_{0}\neq x^{\prime}_{1}$, and define $P$ as before. If w.l.o.g. $\lambda^{\prime}_{0}>\lambda^{\prime}_{1}$, then setting $x_{\beta}:=(1-\beta)x^{\prime}_{0}+\beta x^{\prime}_{1}$, since $F_{|P}=f^{\prime}_{0}=f^{\prime}_{1}$ is affine, we obtain $\lim_{\beta\to+\infty}F(x_{\beta})=-\infty$. So $\lambda^{\prime}_{0}=\lambda^{\prime}_{1}$, i.e. $\mathcal{C}\subset\mathscr{X}\times\{\epsilon\}$ for some $\epsilon\in\mathbb{R}$. Since $0<F$, we necessarily have $\epsilon>0$.
Having shown that $Jf_{\alpha}<\max(Jf_{0},Jf_{1})$ implies (38), as $G$ is bijective, $J^{-1}$ commutes with the max, and as the $p$-coefficient of $J^{-1}(Jf_{\alpha}+\epsilon)$ does not depend on $\epsilon$, we have that $J^{-1}(Jf_{\alpha}+\epsilon)=\langle p_{\alpha},\cdot\rangle+\lambda_{\alpha,\epsilon}$ for some $\lambda_{\alpha,\epsilon}$. We have furthermore that $\lambda_{\alpha,\epsilon}>\lambda_{\alpha}$ since $G_{2}(\_,\cdot)$ is bijective and increasing in $\lambda$. Hence, applying $J^{-1}$ to (38), we derive that
@@
We have just shown that $G$ maps intervals onto intervals. Arguing as in [AAM09, Remark 6, p668] by applying a result of affine geometry, extended in [IRS15, Corollary 3] to arbitrary vector spaces, since $\mathscr{X}\times\mathbb{R}$ has dimension larger than two, $G$ is thus affine and bijective, i.e. we can fix a linear invertible $A:\mathscr{X}^{*}\to\mathscr{X}^{*}$, a linear form $C:\mathscr{X}^{*}\to\mathbb{R}$, and constants $g,b\in\mathscr{X}^{*}$, $d,\delta\in\mathbb{R}$ such that
@@
so, considering nets, we have that $A:\mathscr{X}^{*}\to\mathscr{X}^{*}$ is weak-* continuous as $g(x,\cdot)$ is continuous. Hence $A$ has an adjoint $A^{*}:\mathscr{X}\to\mathscr{X}$ with closed graph. Indeed, by weak-* continuity, for all $x\in\mathscr{X}$, there exists $y\in\mathscr{X}$ such that for all $p\in\mathscr{X}^{*}$, $\langle Ap,x\rangle=\langle p,y_{x}\rangle$. Set $A^{*}(x):=y_{x}$, since $\langle Ap,x\rangle=\langle p,A^{*}(x)\rangle$ and $\mathscr{X}^{*}$ separates the points of the locally convex $\mathscr{X}$, $A^{*}$ is linear. Take a net $(x_{\alpha},y_{\alpha})_{\alpha}$ with $y_{\alpha}=A^{*}x_{\alpha}$ converging to some $(x,y)\in\mathscr{X}\times\mathscr{X}$, then $\langle p,y\rangle=\langle Ap,x\rangle=\langle p,A^{*}x\rangle$, so $y=A^{*}x$ and the graph of $A^{*}$ is closed. Similarly if $(x_{\alpha})_{\alpha}$ converges weakly to some $x\in\mathscr{X}$, then $(A^{*}x_{\alpha})_{\alpha}$ converges weakly to $A^{*}x$.
Fix any convex l.s.c $f$. Then
@@
Conversely, take any $J$ satisfying (37). Such $J$ is clearly order preserving and, reversing the reasoning which led to (39) and using that $A$ has a weak-* continuous adjoint, $Jf$ is the supremum of lower semicontinuous convex functions, hence $Jf\in\mathcal{G}$. If we set $J^{1}f(x)=-\frac{1}{d}\langle b,A^{-1}(x-c)\rangle-\frac{\delta}{d}+\frac{1}{d}\cdot f(A^{-1}(x-c))$, it satisfies $J^{1}Jf=f$.
Applying 11, we know that, for $\mathcal{G}=\operatorname*{Rg}(B)$ for some $\bar{B}$ as in (10), the anti-involutions are only the composition of the reference one $\bar{B}$ with order isomorphisms. In this case, the reference is the Fenchel transform, and we thus recover [AAM09, Theorem 5] for $\mathscr{X}=\mathbb{R}^{d}$ as a special case and derive a straightforward extension to semi-convex functions, i.e. the functions $f$ such that $f+\frac{\alpha}{2}\|\cdot\|^{2}$ is convex l.s.c.
@@
Let $\mathcal{G}$ be the space of proper convex lower semicontinuous functions over a reflexive Banach space $\mathscr{X}$, assumed to be linearly isomorphic to its dual $\mathscr{X}^{*}$. Then the anti-involutions over $\mathcal{G}$ are of the form
@@
with $A\in\mathcal{L}(\mathscr{X})$ invertible, $c\in\mathscr{X}$, $\delta\in\mathbb{R}$ and $K^{-1}A^{-\top}KA=\operatorname*{Id}_{\mathscr{X}}$ where $K:\mathscr{X}\rightarrow\mathscr{X}^{*}$ is the duality operator and $A^{-\top}$ is the inverse of the adjoint of $A$.
@@ Acknowledgments.
The authors deeply thank Sebastián Tapia-García for the fruitful discussions and his many insights on topology. Pierre-Cyril Aubin-Frankowski was funded by the FWF project P 36344-N.
@@
[…]